# PartEdit: Fine-Grained Image Editing using Pre-Trained Diffusion Models

### v1

@@ 4.1 Evaluation Setup
Evaluation Metrics To evaluate the edits, we want to verify that: (1) The edit has been applied at the correct location. (2) The unedited regions and the background have been preserved. Therefore, We evaluate the following metrics for the foreground (the edit) and the background:
  1. 1.
@@
where $\alpha\text{Clip}_{\text{}}$ is the masked CLIP from [33]. There is no overlap between the training images used in token optimization and evaluation (more info in Appendix I).
4.2 Qualitative Results
@@
Synthetic Image Editing Figure 5 shows a qualitative comparison. Our approach excels at performing challenging edits seamlessly while perfectly preserving the background. P2P fails in most cases to perform the edit as it cannot localize the editing part. iP2P completely ignores the specified part and applies the edit to the whole object in some cases and produces distorted images in other cases. Remarkably, our approach succeeds in eliminating racial bias in SDXL by decoupling the hair region from the skin color, as shown in the top-right example.
@@
Real Image Editing Figure 6 shows a qualitative comparison of real image editing. PartEdit produces the most seamless and localized edits, whereas other approaches fail to localize the correct editing region accurately.
4.3 Quantitative Results
@@
Synthetic Image Editing Table 1 top summarizes the quantitative metrics on PartEdit-Synth benchmark. Our approach performs the best on $\alpha\text{Clip}_{\text{avg}}$ and all metrics for the unedited regions. iP2P scores the best in terms $\alpha\text{Clip}_{\text{FG}}$ as it tends to change the whole image to match the editing prompt, ignoring the structure and the style of the original image (see Figure 5). We also performed two user studies comparing our approach against P2P and iP2P on Amazon Mechanical Turk (more details in Appendix H). Our approach is preferred by the users 88.6 % and 77.0% of the time compared to P2P and iP2P, respectively. The reported results for the user studies are based on 360 responses per study.
@@
Mask-Based Editing To demonstrate the effectiveness of our text-based editing approach, we compare it against two mask-based editing approaches, SDXL inpainting and SDXL Latent Blending [2], where the user provides the editing mask. We use the groundtruth annotations to perform the edits for these two approaches, while our approach relies on the optimized tokens. Table 1 shows that our approach is preferred by 75% and 66% of users over the mask-based approaches. This reveals that our editing approach produces visually more appealing edits even compared to the mask-based approaches, where the mask is provided. Our method outperforms and produces more realistic edits compared to both methods, as observed on Figure 8.
@@
Real Image Editing Table 1 shows that PartEdit outperforms Ledits++ and EF-DDPM on all metrics and is significantly favored by users in the user study. This demonstrates the efficacy of our approach in performing fine-grained edits.
@@
SDXL Inpainting [25] 79.72 46.76 34.2 0.96 63.24 75.0
Latent Blending [2] 90.09 46.45 37.5 0.98 68.27 66.1
@@
Ledits++ [3] 66.57 12.95 20.9 0.69 39.76 80.77
EF-DDPM [15] 73.57 15.09 22.9 0.72 44.33 90.38
InfEdit [37] 67.67 21.71 22.5 0.70 44.69 77.88
PnPInversion [16] 71.71 23.59 22.6 0.73 47.65 79.81
@@
Table 1: A quantitative comparison on parts editing. Our Pref. indicates the percentage of users who favored our approach in the user study. Our method outperforms synthetic setting methods without mask (✗), P2P [12] and iP2P [4], and preferred by human preference (Fig. 8) against SDXL inpanting [25], Latent Blending [2] with ground truth masks (✓). Additionally, we showcase the benefits of integrating with existing inversion techniques such as Ledits++ [3].
4.4 Extension to multiple parts
@@
The work focuses mainly on editing a single part at a time. We provide an example of training-free extension to multiple parts, as seen in Figure 7. The extension is done at inference time, with previously separately trained tokens, more details can be seen in Appendix M.
4.5 Ablation Study
@@
Impact of $t_{e}$ The choice of the number of timesteps to perform feature blending using Equation 6 controls the locality of the edit, where the blending always starts at timestep one and ends at $t_{e}$. Figure 10 shows two different scenarios for how to choose $t_{e}$ based on the user desire. In the first row, a low $t_{e}$ causes the horse’s legs and tail to change to dragon ones, but a higher $t_{e}$ would change only the head and preserve everything else. The second row has another example: a lower $t_{e}$ discards the girl’s hair, and a higher $t_{e}$ preserves it. Consequently, $t_{e}$ gives the user more control over the locality of the performed edits.
@@
Impact of Binarization To show the efficacy of our proposed thresholding strategy in Equation 5, we show an edit under different binarization strategies in Figure 11. Standard binarization, where the blending mask is thresholded at 0.5, leads to the least seamless edit as if a robotic mask was placed on the man’s head. OTSU integrates some robotic elements on the neck, but they are not smoothly blended. Finally, our thresholding strategy produces the best edit, where the robotic parts are seamlessly integrated into the neck.
@@
Impact of number or selection of images We further highlight the robustness of token training given the number of images and the selected images in Appendices B and J respectively. Our method outperforms existing mask-based methods (Tab. 1) for part editing under limited data. More ablations in Appendices C and D.
@@
Figure 10: The impact of changing the number of denoising steps $t_{e}$ to perform feature blending.
@@
Figure 11: Comparison of an edit under different mask binarization strategies in our novel layer timestep blending setup.
5 Related Work
5.1 Diffusion-based Image Editing
@@
In general, the image semantics are encoded within the cross-attention layers, which specify where each word in the text prompt is located in the image [12, 34]. In addition, the style and the appearance of the image are encoded through the self-attention layers [35]. Diffusion-based editing approaches exploit these facts to perform different types of semantic or stylistic edits. On the semantic level, several approaches attempt to change the contents of an image according to a user-provided prompt [12, 4, 21, 17, 24]. These changes include swapping an object or altering its surroundings by manipulating the cross-attention maps either through token replacement or attention-map tuning. For stylistic edits, several approaches [5, 35, 24, 13] modify the self-attention maps to apply a specific style to the image while preserving the semantics. This style is either provided by the user in the form of text or a reference image. It is worth mentioning that an orthogonal research direction investigates image inversion to enable real image editing [3, 15, 4]. For a comprehensive review of editing techniques and applications, we refer readers to [14]. Existing text-based editing approaches struggle to apply semantic or stylistic edits to fine-grained object parts, as demonstrated in Section 4. Our approach, PartEdit, is the first step toward enabling fine-grained editing, which will enhance user controllability and experience, and can potentially be integrated into existing editing pipelines.
5.2 Token Optimization
@@
To expand the capabilities of pre-trained diffusion models or to address some of their limitations, they either need to be re-trained or finetuned. However, these approaches are computationally expensive due to the large scale of these models and the training datasets. An attractive alternative is token optimization, where the pre-trained model is kept frozen, and special textual tokens are optimized instead. Those tokens are then used alongside the input prompt to the model to perform a specific task through explicit supervision of cross-attention maps. This has been proven successful in several tasks. [36, 9, 1, 29] employed token optimization and LoRA finetuning to extract or learn new concepts that are used for image generation. [11] learned tokens to detect the most prominent points in images. [22, 18] optimized tokens to perform semantic segmentation. We explored the use of token optimization to learn tokens for object parts customized with a focus on image editing. For this purpose, we adopt SDXL [25] as a base model to obtain the best visual editing quality in contrast to existing approaches that leverage SD 1.5 or 2.1. Our token optimization training is also tailored to obtaining smooth cross-attention maps across all timesteps, unlike [22, 18] that aggregate all attention maps and apply post-processing to obtain binary segmentation masks.
6 Concluding Remarks
@@
Limitations This frozen state also limits our generation’s ability for totally unrealistic edits.
For instance, editing a human head to become a car wheel. Moreover, it can not change the style of the edited part to a different style other than the style of the original image. This limitation stems from the internal design of SDXL that encodes the style in self-attention and prevents mixing two different styles. We provide examples of these scenarios in the Appendix G.
Ethical Impact Our approach can help alleviate the internal racial bias of some edits, such as correlating “Afro” hair with Africans, as we demonstrated in Figure 5. On the other hand, our approach can produce images that might be deemed inappropriate for some users by mixing parts of different animals or humans. Moreover, the remarkable seamless edits performed by our approach can potentially be used for generating fake images, misinformation, or changing the identity.
7 Conclusion
@@ Appendix B Number of Training Samples
Figure 12: The impact of the number of training samples on mIoU.
@@ Appendix C Choice of $\Omega$ for Our Adaptive Thresholding
Figure 13: Applying the edit "with a Joker <head>" with different choices of hyperparameter $\omega$.
@@ Appendix D Choice of Token Padding Strategies
Figure 14: Influence of different token padding strategies during inference on the cross-attention maps.
@@ Appendix E Choice of UNet Layers for Token Optimization
Figure 15: Analysis of how each layer of the UNet performs in terms of mioU. The first eight collected layers of the decoder achieve the best results. Note that we apply OTSU for binarization.
@@ Appendix F Different Edits Per Image
Figure 16: Applying different identity edits to the same image. We showcase the versatility of our method, as there is no change in the underlying model; we can leverage the full capabilities of the model without any retraining or fine-tuning.
@@ Appendix G Failure Cases
Figure 17: Failure cases. Our approach can not perform an edit in a different style (left) or unreasonable edits (right).
@@ Appendix H User study details
Figure 18: Visualization of the user study layout that we conducted.
@@ Appendix I Benchmark and Generating Prompts and Edits for Evaluation
Figure 19: Example of configuration for random generation of the edits.
@@
Figure 20: InfEdit [37] and PnPInversion [16] on real image setting.
Figure 21: An illustration of our user interface.
@@
Figure 22: Additional visualization of obtained attention maps across all time steps of the qualitative results under the real setting.
@@
Segment Anything Kirillov et al. [19], as one of the foundational models in segmentation using conditioned inputs such as points, still struggles to segment parts that do not have a harsh border (commonly torso and head) while it has no problems with classes such as car hood. We can observe such failure cases in Figure 23.
Figure 23: Visualization of masks obtained from Segment Anything (huge) model across the 3 heads for the green provided point. The target indicates what we wanted to segment by the green point.
Appendix P User Interface

### v2

@@ 4.1.4. Evaluation Metrics
To evaluate the edits, we want to verify that: (1) The edit has been applied at the correct location. (2) The unedited regions and the background have been preserved. Therefore, We evaluate the following metrics for the foreground (the edit) and the background:
  1. (1)
@@
where $\alpha\text{Clip}_{\text{}}$ is the masked CLIP from (Sun et al., 2024). There is no overlap between the training images used in token optimization and evaluation (more information in Appendix I).
4.2. Qualitative Results
4.2.1. Synthetic Image Editing
@@
Figure 5 shows a qualitative comparison. Our approach excels at performing challenging edits seamlessly while perfectly preserving the background. P2P fails in most cases to perform the edit as it cannot localize the editing part. iP2P completely ignores the specified part and applies the edit to the whole object in some cases and produces distorted images in other cases. Remarkably, our approach succeeds in eliminating racial bias in SDXL by decoupling the hair region from the skin color, as shown in the top-right example. Additional comparisons in Appendix P .
4.2.2. Real Image Editing
@@
Figure 6 shows a qualitative comparison of real image editing. PartEdit produces the most seamless and localized edits, whereas other approaches fail to localize the correct editing region accurately. Additional comparisons in Appendix L.
4.3. Quantitative Results
4.3.1. Synthetic Image Editing
@@
Table 1 top summarizes the quantitative metrics on PartEdit-Synth benchmark. Our approach performs the best on $\alpha\text{Clip}_{\text{avg}}$ and all metrics for the unedited regions. iP2P scores the best in terms $\alpha\text{Clip}_{\text{FG}}$ as it tends to change the whole image to match the editing prompt, ignoring the structure and the style of the original image (see Figure 5). We also performed two user studies comparing our approach against P2P, iP2P and MasaCtrl on Amazon Mechanical Turk (more details in Appendix H). Our approach is preferred by the users 88.6 %, 77.0% and 73.5% of the time compared to P2P, iP2P and MasaCtrl, respectively. The reported results for the user studies are based on 360 responses per study.
4.3.2. Mask-Based Editing
@@
To demonstrate the effectiveness of our text-based editing approach, we compare it against two mask-based editing approaches, SDXL inpainting and SDXL Latent Blending (Avrahami et al., 2023b), where the user provides the editing mask. We use the groundtruth annotations to perform the edits for these two approaches, while our approach relies on the optimized tokens. Table 1 shows that our approach is preferred by 75% and 66% of users over the mask-based approaches. This reveals that our editing approach produces visually more appealing edits even compared to the mask-based approaches, where the mask is provided. Our method produces more realistic edits compared to both methods, as seen on Figure 10.
4.3.3. Real Image Editing
@@
Table 1 shows that PartEdit outperforms Ledits++, EF-DDPM and other methods on all metrics and is significantly favored by users in the user study. This demonstrates the efficacy of our approach in performing fine-grained edits.
Table 1. A quantitative comparison on parts editing. Our Pref. indicates the % of users who favored our approach in the user study. Our method outperforms synthetic setting methods without mask (✗) (P2P (Hertz et al., 2022), iP2P (Brooks et al., 2023), MasaCtrl (Cao et al., 2023)), and preferred by human preference (fig. 10) against SDXL inpanting (Podell et al., 2024), Latent Blending (Avrahami et al., 2023b) with ground truth masks (✓). Additionally, we showcase the benefits of integrating with existing inversion techniques such as Ledits++ (Brack et al., 2024).
@@
SDXL Inpainting (Podell et al., 2024) 79.72 46.76 34.2 0.96 63.24 75.0
Latent Blending (Avrahami et al., 2023b) 90.09 46.45 37.5 0.98 68.27 66.1
@@
Ledits++ (Brack et al., 2024) 66.57 12.95 20.9 0.69 39.76 80.77
EF-DDPM (Huberman-Spiegelglas et al., 2024) 73.57 15.09 22.9 0.72 44.33 90.38
InfEdit (Xu et al., 2024) 67.67 21.71 22.5 0.70 44.69 77.88
PnPInversion (Ju et al., 2024b) 71.71 23.59 22.6 0.73 47.65 79.81
TurboEdit (Deutch et al., 2024) 75.55 12.49 20.2 0.69 44.02 66.92
ReNoise (Garibi et al., 2024) 69.80 9.93 20.7 0.68 39.86 77.69
@@
4.4. Extension to multiple parts
@@
The work focuses mainly on editing a single part at a time. We provide an example of training-free extension to multiple parts, as seen in Figure 7. The inference time extension with previously separately trained tokens, more details can be seen in Appendix M.
4.5. Ablation Study
4.5.1. Impact of $t_{e}$
@@
The choice of the number of timesteps to perform feature blending using Equation 6 controls the locality of the edit, where the blending always starts at timestep one and ends at $t_{e}$. Figure 8 shows two different scenarios for how to choose $t_{e}$ based on the user desire. In the first row, a low $t_{e}$ causes the horse’s legs and tail to change to dragon ones, but a higher $t_{e}$ would change only the head and preserve everything else. The second row has another example: a lower $t_{e}$ discards the girl’s hair, and a higher $t_{e}$ preserves it. Consequently, $t_{e}$ gives the user more control over the locality of the performed edits.
4.5.2. Impact of Binarization
@@
To show the efficacy of our proposed thresholding strategy in Equation 5, we show an edit under different binarization strategies in Figure 9. Standard binarization, where the blending mask is thresholded at 0.5, leads to the least seamless edit as if a robotic mask was placed on the man’s head. OTSU integrates some robotic elements on the neck, but they are not smoothly blended. Finally, our thresholding strategy produces the best edit, where the robotic parts are seamlessly integrated into the neck.
4.5.3. Impact of number or selection of images
@@
We further highlight the robustness of token training given the number of images and the selected images in Appendices B and J respectively. Our method outperforms existing mask-based methods (table 1) for part editing under limited data. In appendix P, we ablate off-the-shelf models (i.e. Grounded SAM (Ren et al., 2024) using Kirillov et al. (2023); Liu et al. (2023)). More ablations in Appendices C and D.
@@
Figure 8. The impact of the number of denoising steps $t_{e}$ to perform feature blending.
@@
Figure 9. Comparison of an edit under different mask binarization strategies in our novel layer timestep blending setup.
Figure 10. A comparison on synthetic benchmark against Latent Blending (Avrahami et al., 2023b) and SDXL Inpainting (Podell et al., 2024) using ground truth masks (✓) against our predicted masks (✗). We observe PartEdit outperforming Latent Blending, which fails totally in some of the edits (spiderman, hair, and chair), while others produce unintegrated edits, such as the bear head. More detailed comparisons can be found on the website.
Figure 11. Different edits per image on real image editing using our method. We showcase the versatility of our method, as there is no change in the underlying model; we can leverage the full capabilities of the model without any retraining or fine-tuning. More details in Appendix F .
Figure 12. Challenging multiple subjects edits (more in fig. 26 ). Examples of Grounded SAM (Ren et al., 2024) integration for "second dog" or "right" respectively.
5. Related Work
5.1. Diffusion-based Image Editing
@@
In general, the image semantics are encoded within the cross-attention layers, which specify where each word in the text prompt is located in the image (Hertz et al., 2022; Tang et al., 2023). In addition, the style and the appearance of the image are encoded through the self-attention layers (Tumanyan et al., 2023). Diffusion-based editing approaches exploit these facts to perform different types of semantic or stylistic edits. On the semantic level, several approaches attempt to change the contents of an image according to a user-provided prompt (Hertz et al., 2022; Brooks et al., 2023; Lin et al., 2023; Kawar et al., 2023; Parmar et al., 2023; Cao et al., 2023; Avrahami et al., 2024). These changes include swapping an object or altering its surroundings by manipulating the cross-attention maps either through token replacement or attention-map tuning. For stylistic edits, the self-attention maps are commonly modified by several approaches (Cao et al., 2023; Tumanyan et al., 2023; Parmar et al., 2023; Hertz et al., 2023) to apply a specific style to the image while preserving the semantics. This style is either provided by the user in the form of text or a reference image. It is worth mentioning that an orthogonal research direction investigates image inversion to enable real image editing (Brack et al., 2024; Huberman-Spiegelglas et al., 2024; Brooks et al., 2023; Deutch et al., 2024; Garibi et al., 2024). For a comprehensive review of editing techniques and applications, we refer readers to (Huang et al., 2024). Existing text-based editing approaches struggle to apply semantic or stylistic edits to fine-grained object parts, as demonstrated in Section 4. Our approach, PartEdit, is the first step toward enabling fine-grained editing, which will enhance user controllability and experience, and can potentially be integrated into existing editing pipelines.
5.2. Token Optimization
@@
To expand the capabilities of pre-trained diffusion models or to address some of their limitations, they either need to be re-trained or finetuned. However, these approaches are computationally expensive due to the large scale of these models and the training datasets. An attractive alternative is token optimization, where the pre-trained model is kept frozen, and special textual tokens are optimized instead. Those tokens are then used alongside the input prompt to the model to perform a specific task through explicit supervision of cross-attention maps. This has been proven successful in several tasks. (Valevski et al., 2023; Gal et al., 2022; Avrahami et al., 2023a; Safaee et al., 2024) employed token optimization and LoRA finetuning to extract or learn new concepts that are used for image generation. (Hedlin et al., 2023) learned tokens to detect the most prominent points in images. (Marcos-Manchón et al., 2024; Khani et al., 2024) optimized tokens to perform semantic segmentation. We explored the use of token optimization to learn tokens for object parts customized with a focus on image editing. For this purpose, we adopt SDXL (Podell et al., 2024) as a base model to obtain the best visual editing quality in contrast to existing approaches that leverage SD 1.5 or 2.1. Our token optimization training is also tailored to obtaining smooth cross-attention maps across all timesteps, unlike (Marcos-Manchón et al., 2024; Khani et al., 2024) that aggregate all attention maps and apply post-processing to obtain binary segmentation masks.
6. Concluding Remarks
@@
Limitations
This frozen state also limits our generation’s ability for totally unrealistic edits.
For instance, editing a human head to become a car wheel. Moreover, it can not change the style of the edited part to a different style other than the style of the original image. This limitation stems from the internal design of SDXL that encodes the style in self-attention and prevents mixing two different styles. We provide examples of these scenarios in the Appendix G . Moreover, we provide analysis of Flux (Black Forest Labs, 2024), a DiT-based model, which shows promising improvements in base model localization in Appendix R.
Ethical Impact
Our approach can help alleviate the internal racial bias of some edits, such as correlating “Afro” hair with Africans, as we demonstrated in Figure 5. On the other hand, our approach can produce images that might be deemed inappropriate for some users by mixing parts of different animals or humans. Moreover, the remarkable seamless edits performed by our approach can potentially be used for generating fake images and misinformation.
7. Conclusion
@@
Acknowledgements.
We thank the anonymous reviewers for their constructive comments. The work is supported by funding from KAUST - Center of Excellence for Generative AI, under award number 5940, and the NTGC-AI program.
@@ Appendix B Number of Training Samples
Figure 13. The impact of the number of training samples on mIoU.
@@ Appendix C Choice of $\Omega$ for Our Adaptive Thresholding
Figure 14. Applying the edit "with a Joker <head>" with different choices of hyperparameter $\omega$.
@@ Appendix D Choice of Token Padding Strategies
Figure 15. Influence of different token padding strategies during inference on the cross-attention maps.
@@ Appendix E Choice of UNet Layers for Token Optimization
Figure 16. Analysis of how each layer of the UNet performs in terms of mioU. The first eight collected layers of the decoder achieve the best results. Note that we apply OTSU for binarization.
@@ Appendix F Different Edits Per Image
Figure 17. Applying different identity edits to the same image. We showcase the versatility of our method, as there is no change in the underlying model; we can leverage the full capabilities of the model without any retraining or fine-tuning.
@@ Appendix G Failure Cases
Figure 18. Failure cases. Our approach can not perform an edit in a different style (left) or unreasonable edits (right).
@@ Appendix H User study details
Figure 19. Visualization of the user study layout that we conducted.
@@ Appendix I Benchmark and Generating Prompts and Edits for Evaluation
Figure 20. Example of configuration for random generation of the edits.
@@
In Figure 24, we present a comparison with ReNoise and TurboEdit on the same examples. ReNoise fails to apply the edit in four cases, exhibits inversion failure in one case, and successfully edits the "alien head" example, though with poor background preservation. In contrast, TurboEdit performs three edits, one of which changes identity (curly hair). When edits are unsuccessful, it mostly preserves the original image, with only minor alterations.
Figure 21. InfEdit (Xu et al., 2024) and PnPInversion (Ju et al., 2024b) on real image setting.
Figure 22. Qualitative comparison on synthetic images from the PartEdit benchmark, evaluating Grounded SAM (Ren et al., 2024) (in our setting, with post-processing as described in Appendix P) and the additional baseline MasaCtrl (Cao et al., 2023). MasaCtrl fails to perform the edit in most cases and alters identity in the hair example. Using Grounded SAM in our mask component achieves comparable results in 3 edits, and showcases a trend of localizing to the whole object (e.g. car and chair). We explore more of this phenomenon in Figure 27.
Figure 23. Visualization for the cross-attention maps of FLUX (Black Forest Labs, 2024) that corresponds to different words of the textual prompt. DiT-based models show greater promise for accurate localization. The first example shows that the whole head is not covered, while the second interprets the "black hood" as a "black stripe."
Figure 24. Comparison against ReNoise (Garibi et al., 2024) and TurboEdit (Deutch et al., 2024) on real image setting.
Figure 25. DiT-based editing results in a synthetic setting using StableFlow (Avrahami et al., 2024). A clear gap can be observed between generation (fig. 23) and editing performance of the FLUX model shown above. In the first row, the edited part is localized to a smaller region than expected. In the second row, the edits are misaligned; either applied to the wrong location (e.g., destroyed car hood), appear layered on top of the original object, or affect the entire car instead of the intended part.
Figure 26. Additional challenging examples of multiple subject edits.
Figure 27. Visualization of the generated image, annotated ground truth, and segmentation masks obtained using Grounded SAM (Ren et al., 2024), which combines Grounding DINO (Liu et al., 2023) and SAM (Kirillov et al., 2023). We also show the average attention maps across all timesteps in our setting (see appendix P). Grounded SAM struggles to segment fine-grained parts such as "chair seat," "torso," "car body," and "car hood," often returning masks that encompass the entire object.
Table 2. Quantitative metrics of using off-the-shelf Grounded SAM (Ren et al., 2024) masks. We use our setting, with details described in Appendix P.
Method Edited Region Unedited Region $\alpha\text{Clip}_{\text{avg}}$$\uparrow$
$\alpha\text{Clip}_{\text{FG}}$$\uparrow$ $\alpha\text{Clip}_{\text{BG}}$$\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
PartEdit + GroundedSAM masks 92.37 23.61 28.2 0.92 57.99
PartEdit + Our token masks 91.74 38.38 31.5 0.98 65.06
Figure 28. An illustration of our user interface.
@@
Figure 29. Additional visualization of obtained attention maps across all time steps of the qualitative results under the real setting.
@@
Segment Anything Kirillov et al. (2023), as one of the foundational models in segmentation using conditioned inputs such as points, still struggles to segment parts that do not have a harsh border (commonly torso and head) while it has no problems with classes such as car hood. We can observe such failure cases in Figure 30.
Figure 30. Visualization of masks obtained from Segment Anything (huge) model across the 3 heads for the green provided point. The target indicates what we wanted to segment by the green point.
Appendix P Comparison against Grounded-SAM and MasaCtrl
We provide a comparison and analysis against the off-the-shelf segmentation model Grounded-SAM (Ren et al., 2024) in our synthetic setting. We utilize binary masks from Grounded-SAM in the mask component of the pipeline shown in Figure 3. For both the DINO (Liu et al., 2023) and SAM (Kirillov et al., 2023) models in GroundedSAM, we use the base models with the HuggingFace Transformers (Wolf et al., 2020) implementation [^3] https://github.com/NielsRogge/Transformers-Tutorials/tree/master/Grounding%20DINO.
Since SAM produces multiple overlapping masks, we first select the smallest mask from each group of overlapping masks. Then, when multiple instances are detected, we choose the largest of these smallest masks to represent the main object of interest. We use a threshold of 0.3 with polygonal refinement for SAM, and a threshold of 0.1 for overlap during mask selection.
Figure 27 shows a visualization of masks predicted by Grounded-SAM and our optimized tokens, as well as the groundtruth, The figure shows that Grounded-SAM often fails to segment parts and segments the whole object instead. This aligns with the observation made in Figure 30 where SAM model favors object-level segmentation masks which makes it challenging to robustly segment parts. We also report a quantitative comparison in Table 2, where foreground metrics for the edited region are higher because of this inherent preference for segmenting object. On the other hand, the background metrics are worse as the whole object is edited rather than part. We also provide some editing examples in Figure 22.
We also provide a qualitative comparison for MasaCtrl (Cao et al., 2023) in Figure 22. The figure shows that MasaCtrl performs poorly of fine-grained editing.
Appendix Q Token optimization duration
The optimization time depends on the number of optimized layers, model size, and training set size. When optimizing the first 8 layers of the decoder (optimal setup as shown in Figure 16), the optimization takes 330 and 129 seconds for SDXL and SD 2.1, respectively, with 10 images and 1000 optimization steps using A100 in FP32 precision.
Appendix R Analysis of DiT-based models
To assess whether the failure cases we observe are specific to U-Net-based architectures, we conducted a supplemental experiment using a transformer-based diffusion model, Flux (Black Forest Labs, 2024). We employed an existing attention visualization implementation [^4] https://github.com/wooyeolbaek/attention-map-diffusers, running $1024^{2}$ inference of Flux.dev with 30 inference steps, a guidance scale of 3.5, and seed 420. Figure 23 shows an improved localization and better adherence to the prompt as a result of employing a full transformer-based architecture.
Additionally, we use StableFlow (Avrahami et al., 2024), a recent synthetic editing work that focuses on DiT-based models, to check editing capabilities. In Figure 25, we can observe a localization performance gap with often artifacts or poor localization compared to direct generation in Figure 23.
Appendix S User Interface